OpenStudy (anonymous):

Find the exact value of the hypotenuse of the triangle. Use the triangle below to answer this question.

OpenStudy (anonymous):



OpenStudy (anonymous):

cos 30 = 9/x

OpenStudy (anonymous):

Use Cos function

OpenStudy (anonymous):

@CynthiaJune94 I think you're too much confused in Trigonometry

OpenStudy (anonymous):

I am!! I dont know how to do it.... :( and would that be 270?

OpenStudy (anonymous):

no it wouldn't

OpenStudy (anonymous):

then how would you solve that? wouldnt you multiply the nine on both sides to get the x by itself?

OpenStudy (anonymous):

if you multiplied 9 on both sides you would get: 9 cos 30 = 81/x

OpenStudy (anonymous):

okay then how do you do it right? would 7.794 be the correct answer?

OpenStudy (anonymous):

how did you get 7.794?

OpenStudy (anonymous):

cos 30=.866 then i multiplied that by 9 oh maybe i should have divided....

OpenStudy (anonymous):

where did you get .866?

OpenStudy (anonymous):

i did the cos button on my graphing calculator and multplied that by 30

OpenStudy (anonymous):

you don't multiply cos by 30, cos is an operation. you're "cosing" 30. cos 30 is one term.

OpenStudy (anonymous):

yea it looked like this on my calculator... cos(30)=.866

OpenStudy (anonymous):

rearrange and solve for x

OpenStudy (anonymous):

.866 = 9/x

OpenStudy (anonymous):

learn how to do it without a graphing calculator, you won't be able to use it on a test.

OpenStudy (anonymous):

x=7.794 Right?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

sigh.... i dont know....

OpenStudy (anonymous):

im just not getting it...

OpenStudy (anonymous):

3 = 9/x whats x?

OpenStudy (anonymous):

x=27 where did you get 3 tho?

OpenStudy (anonymous):

i'm giving you an easier example.

OpenStudy (anonymous):

so you think 9 / 27 = 3?

OpenStudy (anonymous):

no....

OpenStudy (anonymous):

9/x=.866?

OpenStudy (anonymous):

yes, so what's x?

OpenStudy (anonymous):

you learned this in 7th grade. if 3 = 9/x what do you do to solve for x?

OpenStudy (anonymous):

so x would be 10.345

OpenStudy (anonymous):

you multiply both sides by x and then divide both sides by 3 to make x the subject of your equations. with that you get 3

OpenStudy (anonymous):

x=10.3926

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

how did you get that? yes

OpenStudy (anonymous):

he said multiply both sides by x making it .866x=9 then i divided 9 by .866 x=10.3926 Omg!!! Thank you so much but i have one more like this i need help with

OpenStudy (anonymous):

cynthia, i hope you've understood the concept

OpenStudy (anonymous):

see my next question and see if i did it right? PLZ